# The synthetic instrument: From sparse association to sparse causation

### v2

@@ 3.2 The synthetic instrument
On the surface, one may follow the identification strategy described in Section 3.1 to estimate $\beta$. However, in the general case where $q>1$, each expert would hypothesize that exactly $q$ elements of $\beta$ are zero. In total, there are $C_{p}^{q}$ different hypotheses. Several challenges arise when the data is moderate to high-dimensional, so that $p$ and $q$ are not small.
@@
Other approaches that use the voting analogy for identification (e.g. Zhou et al., 2014; Guo et al., 2018) face the same challenges we present here. It is only due to the special structure of our problem that we are able to develop a method that bypasses the model selection step and addresses these challenges. See Sun et al. (2023) for another example in a different context.
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are identified up to the same sign flip so that $\Lambda_{1}/\Lambda_{2}$ is identifiable. Eliminating $U$ from (5), we get $X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}=\epsilon_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}\epsilon_{2}$, which only depends on the error terms $\epsilon_{1}$ and $\epsilon_{2}$. Since $\epsilon_{2}$ is also uncorrelated with $U$, it is not difficult to see from Figure 1(b) that $SIV_{1}^{(2)}=X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}$ satisfies the conditions for an instrumental variable for identifying $\beta_{1}$ described in Section 2.3, hence the name synthetic instrument. In contrast to a standard instrumental variable, the synthetic instrument is directly constructed as a linear combination of the exposures, so there is no need to measure additional exogenous variables.
To identify $\beta_{3}$, one can similarly define $SIV_{1}^{(3)}=X_{3}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{3}.$ Let $SIV^{(2)}=\left(SIV_{1}^{(2)},SIV_{3}^{(2)}\right)$. One can then obtain $(\beta_{1},\beta_{3})$ using the so-called synthetic two-stage least squares:
@@
    Fit a linear regression of $X=(X_{1},X_{2},X_{3})$ on $SIV^{(2)}=(SIV_{1}^{(2)},SIV_{3}^{(2)})$ and obtain $\widetilde{X}=E[X\mid SIV^{(2)}]$.
@@ 3.3 Voting with the synthetic instrument
    For $j=1,2,3$, fit a linear regression of $X$ on $SIV^{(j)}$ and obtain $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$;
@@
    Fit a linear regression of $Y$ on $\widetilde{X}^{(j)}$, fixing $\beta_{j}=0$, and obtain the coefficients $\widetilde{\beta}^{(j)}$.
@@
On the surface, similar to the problems described at the beginning of Section 3.2, voting with the synthetic instrument still involves fitting three different regressions and comparing three vectors $\widetilde{\beta}^{(j)}.$ We now make two key observations regarding the properties of the synthetic instrument, which allows us to simplify Algorithm 1 into a two-stage regression procedure.
@@
Let $\Lambda=(\Lambda_{1},\Lambda_{2},\Lambda_{3})$. For $j=1,2,3$, $SIV^{(j)}\in\mathbb{R}^{2}$ span the same linear space $\{\lambda^{T}X:\lambda\in\Lambda^{\perp}\}.$ As a result, $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$ does not depend on the choice of $j$ so that one only needs to run Step 1 of Algorithm 1 once.
@@
From Table 1, we observe that $\|\widetilde{\beta}^{(1)}\|_{0}=\|\widetilde{\beta}^{(2)}\|_{0}=1$, while $\|\widetilde{\beta}^{(3)}\|_{0}=2$. Recall that the true value is $\dot{\beta}=\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}$. Instead of calculating $\widetilde{\beta}^{(j)},j=1,2,3$ separately for each $j$, Steps 2 and 3 in Algorithm 1 can be replaced with the following penalized regression:
$\beta^{SIV}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{3}}\|\ Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\|_{2}^{2}\text{\quad subject to }\|\beta\|_{0}\leq 1,$
@@ 3.4 Synthetic two-stage regularized regression
We now formally introduce the synthetic two-stage regularized regression for the general case. Motivated by Observation 1, we introduce the following definition for the synthetic instrument.
@@
Under models (1), (3), and Condition A2, the $SIV$ given by Definition 1 serve as valid instrumental variables for estimating the treatment effects of $X$ on $Y$.
@@
    (Plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$, and suppose that $\dot{\beta}_{C^{*}}\neq 0$. The synthetic two-stage least squares coefficient obtained by assuming $\beta_{C^{*}}=0$ is given by $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}\,\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}$, where $\widetilde{X}=\mathbb{E}(X\mid SIV)$. The plurality rule assumes that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q$.
In Condition A4, each $C^{*}$ corresponds to an expert who makes the incorrect hypothesis that $\beta_{C^{*}}=0$. Let $s=\|\dot{\beta}\|_{0}$. In general, there are $C_{p-s}^{q}$ experts making correct hypotheses. If $s<p-q$, then there are at least $q+1$ experts making correct hypotheses. The plurality rule assumes that no more than $q$ incorrect hypotheses lead to the same synthetic two-stage least squares coefficient. This assumption is similar in spirit to the plurality assumption used in the invalid IV literature (e.g. Guo et al., 2018). We discussed this assumption in more detail in Section LABEL:sec:disA4 of the supplementary material.
@@
Suppose that models (1), (3) and conditions A1, A2, and A4 hold.
@@
An important feature of Theorem 1 is that given $q$, it is possible to test the sparsity condition A3 from the observed data. In particular, it shows that under models (1), (3), conditions A1, A2, and the plurality rule A4, the following three statements are equivalent:
@@
    The most sparse least-squares solution to the second-stage regression has an $\ell_{0}$-norm smaller than $p-q$, i.e.,
@@
It is worth noting that (6) can be checked from the observed data distribution, so that one may develop a consistent test for Condition A3 and identifiability of $\beta$ under models (1), (3) and conditions A1, A2, and A4. See Algorithm 2 below for more details.
@@ 4.1 Estimation
where $k$ is a tuning parameter, $\widehat{\Lambda}$ is an estimator of the loading matrix, and $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}.$
Several different estimators have been proposed for determining the number of latent factors in a factor model. In our simulations and data analysis, we will use the estimator developed by Onatski (2010) to obtain $\widehat{q}$, as it is applicable in both low- and high-dimensional settings. Similarly, numerous estimators have been developed for the loading matrix $\Lambda$. For low-dimensional settings where $p$ is fixed, we recommend using the maximum likelihood estimator for $\Lambda$ by optimizing the log-likelihood under the assumption of multivariate normality. For high-dimensional settings, we suggest estimating $\Lambda$ using the principal component analysis (PCA) method (Bai, 2003), which provides a consistent estimator for the loading matrix without assuming the covariance matrix $\text{Cov}(\epsilon_{X})$ to be diagonal. Finally, we use cross-validation to select the tuning parameter $k$. Algorithm 2 summarizes our estimation procedure.
@@
1:Obtain $\widehat{q}$ from $\bm{X}$ (e.g., Onatski, 2010).
@@
5:Obtain $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}$.
6:Obtain $\widehat{\bf X}=\widehat{\mathbb{E}}({\bf{X}}\mid\widehat{SIV})$ via ordinary least squares.
@@ 4.2 Theoretical properties
In this section, we study the theoretical properties of the estimator $\widehat{\beta}$ in Algorithm 2. We consider two paradigms: (1) low-dimensional settings where the dimension of exposure, $p$, is fixed; and (2) high-dimensional settings where $p$ grows with the sample size $n$. For the former, we show that under mild regularity conditions, $\widehat{\beta}$ is $\sqrt{n}$-consistent. For the latter, we show that under mild regularity conditions, $\widehat{\beta}$ achieves a lasso-type error bound. We also demonstrate variable selection consistency in both scenarios. In our theoretical results, we do not require $\widehat{\Lambda}=\Lambda$. We only need certain norms of $\widehat{\Lambda}$ to be consistent with $\Lambda$, which can be achieved by classical estimators.
We first introduce assumptions for the low-dimensional case.
@@
    All coefficients ${\Lambda}$, $\beta$, and the function $g(\cdot)$ in models (1) and (3) are fixed and do not change as $n$ goes to infinity.
@@
    $U_{i}$, $\epsilon_{x,i}$, and $\epsilon_{y,i}$ are independent random draws from the joint distribution of $(U,\epsilon_{x},\epsilon_{y})$ such that $E(\epsilon_{x})=\bm{0}$, $E(U)=\bm{0}$, $\text{Cov}(\epsilon_{x})=D$, $\text{Cov}(U)=I_{q}$, and $(U,\epsilon_{x},\epsilon_{y})$ are mutually independent. Furthermore, assume that $\text{Var}({\epsilon}_{y})={\sigma}^{2}$ and $\max_{1\leq j\leq p}\text{Var}(X_{j})=\sigma_{x}^{2}$; these parameters are fixed and do not change as $n$ goes to infinity.
@@
    Let ${\Sigma}_{\widetilde{X}}=\text{Cov}(\widetilde{X})$. We assume $\min_{\theta\in\mathbb{R}^{p},0<\|\theta\|_{0}\leq 2s}\frac{\theta^{\mathrm{\scriptscriptstyle T}}\Sigma_{\widetilde{X}}\theta}{\|\theta\|_{2}^{2}}>c$ for some positive constant $c$.
Conditions B1B2 are standard assumptions for the low-dimensional setting. Given Condition A2, Condition B3 assumes that the estimator for factor loading is root-$n$ consistent. Condition B4 is the population version of the sparse eigenvalue condition (Raskutti et al., 2011, Assumption 3(b)).
@@
In Theorem 2, it is assumed that $\widehat{k}=s$. This is a standard condition in the $\ell_{0}$-optimization literature (e.g. Raskutti et al., 2011; Shen et al., 2013).
@@
    $sq^{2}\log(p)\log(n)/n\rightarrow 0$, $n=O(p)$, and $q+\log(p)\lesssim\sqrt{n}$, where $x\lesssim y$ means there exist a constant $C$ such that $x\leq Cy$.
@@
    Assume the random variables in models (1) and (3) satisfy $E(X)=\bm{0}$, $E(U)=\bm{0}$, and $\text{Cov}(U)=I_{q}$. We also assume $\epsilon_{y}$ is independent of $(X,U)$ and ${\epsilon_{x}}$ is independent of $U$. Furthermore, assume ${\epsilon}_{y}$, $\epsilon_{x,j}$, and $X_{j}$ are sub-Gaussian random variables with sub-Gaussian parameters ${\sigma}^{2}$, $\widetilde{\sigma}_{j}^{2}$, and $\sigma_{j}^{2}$. The parameters satisfy ${\sigma}^{2}\leq C_{4}$, and $C_{5}\leq\widetilde{\sigma}_{j}^{2},\sigma_{j}^{2}\leq C_{6}$ for some constants $C_{4},C_{5},C_{6}>0$.
@@
Condition C1 allows the number of exposures $p$ to grow exponentially with the sample size, while the number of latent confounders $q$ grows at a slower polynomial rate. Condition C2 is a standard assumption in high-dimensional factor analysis (Fan et al., 2013; Shen et al., 2016) for loading identification. Condition C3 assumes that the exposures $X_{j}$ are sub-Gaussian, and the noise level is bounded. Condition C4 is a standard assumption on minimum signal strength.
@@
Assuming that Conditions A1A4 and C1C3 hold, and the tuning parameter $\widehat{k}=s$, we have:
@@
where $b$ is the best linear predictor of $U$ from $X$. Their result depends on (i) $\|b\|_{2}=O_{p}\left(\frac{1}{\sqrt{p}}\right)$ and (ii) the term $(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y}$ being independent of $X$ under their joint Gaussian assumption. However, these assumptions may not hold in our model (3), where $g(U)$ is an unknown function.

### v3

@@ 3.2 The synthetic instrument
On the surface, one may follow the identification strategy described in Section 3.1 to estimate $\beta$. However, in the general case where $q>1$, each expert would hypothesize that exactly $q$ elements of $\beta$ are zero. In total, there are $C_{p}^{q}$ different hypotheses. Several challenges arise when the data are moderate to high-dimensional, so that $p$ and $q$ are not small.
@@
Other approaches that use the voting analogy for identification (e.g., Zhou et al., 2014; Guo et al., 2018) face the same challenges we present here. It is only due to the special structure of our problem that we are able to develop a method that bypasses the model selection step and addresses these challenges.
@@
where $\Lambda_{1}$ and $\Lambda_{2}$ are identified up to the same sign flip, so that $\Lambda_{1}/\Lambda_{2}$ is identifiable. Eliminating $U$ from (5), we obtain $X_{1}-{\Lambda_{2}}X_{2}/{\Lambda_{1}}=\epsilon_{1}-{\Lambda_{1}}\epsilon_{2}/{\Lambda_{1}},$ which depends only on the error terms $\epsilon_{1}$ and $\epsilon_{2}$. Since $\epsilon_{2}$ is also uncorrelated with $U$, it is not difficult to see from Figure 1(b) that $SIV_{1}^{(2)}=X_{1}-{\Lambda_{1}}X_{2}/{\Lambda_{2}}$ satisfies the conditions for an instrumental variable for identifying $\beta_{1}$ described in Section 2.3, hence the name synthetic instrument. In contrast to a standard instrumental variable, the synthetic instrument is directly constructed as a linear combination of the exposures, so there is no need to measure additional exogenous variables.
To identify $\beta_{3}$, one can similarly define $SIV_{3}^{(2)}=X_{3}-{\Lambda_{3}}X_{2}/{\Lambda_{2}}$. Let $SIV^{(2)}=\left(SIV_{1}^{(2)},SIV_{3}^{(2)}\right)$. One can then obtain $(\beta_{1},\beta_{3})$ using the so-called synthetic two-stage least squares:
@@
    Fit a linear regression of $X=(X_{1},X_{2},X_{3})$ on $SIV^{(2)}=(SIV_{1}^{(2)},SIV_{3}^{(2)})$ and obtain $\widetilde{X}=\widehat{\mathbb{E}}[X\mid SIV^{(2)}]$ through the fitted values of the linear regression.
@@ 3.3 Voting with the synthetic instrument
    For $j=1,2,3$, fit a linear regression of $X$ on $SIV^{(j)}$ and obtain $\widetilde{X}^{(j)}=\widehat{\mathbb{E}}\left[X\mid SIV^{(j)}\right]$;
@@
    Fit a linear regression of $Y$ on $\widetilde{X}^{(j)}$, fixing $\beta_{j}=0$, and obtain the coefficients $\widetilde{\beta}^{(j)}$;
@@
On the surface, similar to the problems described at the beginning of Section 3.2, voting with the synthetic instrument still involves fitting three different regressions and comparing three vectors $\widetilde{\beta}^{(j)}$. We now make two key observations regarding the properties of the synthetic instrument, which allow us to simplify Algorithm 1 into a two-stage regression procedure.
@@
Let $\Lambda=(\Lambda_{1},\Lambda_{2},\Lambda_{3})$. For $j=1,2,3$, $SIV^{(j)}\in\mathbb{R}^{2}$ span the same linear space $\{\lambda^{\mathrm{\scriptscriptstyle T}}X:\lambda\in\Lambda^{\perp}\}$. As a result, $\widetilde{X}^{(j)}=\mathbb{E}\left[X\mid SIV^{(j)}\right]$ does not depend on the choice of $j$, so that one only needs to run Step 1 of Algorithm 1 once.
@@
From Table 1, we observe that $\|\widetilde{\beta}^{(1)}\|_{0}=\|\widetilde{\beta}^{(2)}\|_{0}=1$, while $\|\widetilde{\beta}^{(3)}\|_{0}=2$. Recall that the true value is $\dot{\beta}=\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}$. Instead of calculating $\widetilde{\beta}^{(j)},\,j=1,2,3,$ separately for each $j$, Steps 2 and 3 in Algorithm 1 can be replaced with the following penalized regression:
$\beta^{SIV}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{3}}\|Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\|_{2}^{2}\quad\text{subject to }\|\beta\|_{0}\leq 1,$
@@ 3.4 Synthetic two-stage regularized regression
We now formally introduce the synthetic two-stage regularized regression for the general case. Motivated by Observation 1, we provide the following definition of the synthetic instrument.
@@
Under models (1), (3), and Condition A2, the SIVSIV given by Definition 1 serve as valid instrumental variables for estimating the treatment effects of $X$ on $Y$.
@@
    (Plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$, and suppose that $\dot{\beta}_{C^{*}}\neq 0$. The synthetic two-stage least squares coefficient obtained by assuming $\beta_{C^{*}}=0$ is given by $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}\,\mathbb{E}\bigl(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\bigr)^{2},$ where $\widetilde{X}=\widehat{\mathbb{E}}(X\mid SIV)$. The plurality rule assumes that $\max\limits_{\beta\in\mathbb{R}^{p}}\,|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
In Condition A4, each $C^{*}$ corresponds to an expert who makes the incorrect hypothesis that $\beta_{C^{*}}=0$. Let $s=\|\dot{\beta}\|_{0}$. In general, there are $C_{p-s}^{q}$ experts making correct hypotheses. If s<pqs<p-q, then there are at least $q+1$ experts making correct hypotheses. The plurality rule assumes that no more than $q$ incorrect hypotheses lead to the same synthetic two-stage least squares coefficient. This assumption is similar in spirit to the plurality assumption used in the invalid IV literature (e.g., Guo et al., 2018). We further discuss Assumption A4 in Section LABEL:sec:disA4 of the supplementary material, where we argue that its violation is unlikely.
@@
Suppose that models (1), (3), and Conditions A1, A2, and A4 hold.
@@
An important feature of Theorem 1 is that, given $q$, it is possible to test the sparsity condition A3 from the observed data. In particular, it shows that under models (1), (3), Conditions A1, A2, and the plurality rule A4, the following three statements are equivalent:
@@
    The most sparse least-squares solution to the second-stage regression has an $\ell_{0}$-norm smaller than pqp-q, i.e.,
@@
It is worth noting that (6) can be checked from the observed data distribution, so that one may develop a consistent test for Condition A3 and the identifiability of $\beta$ under models (1), (3), and Conditions A1, A2, and A4. See Algorithm 2 below for more details.
@@ 4.1 Estimation
where $k$ is a tuning parameter, $\widehat{\Lambda}$ is an estimator of the loading matrix, and $\widehat{SIV}={\bf X}B_{\widehat{\Lambda}^{\perp}}.$
Several estimators have been proposed to determine the number of latent factors in a factor model. In our simulations and data analysis, we use the estimator developed by Onatski (2010) to obtain $\widehat{q}$, as it is applicable in both low- and high-dimensional settings. Likewise, various methods exist for estimating the loading matrix $\Lambda$. In low-dimensional settings where $p$ is fixed, we recommend the maximum likelihood estimator of $\Lambda$, obtained by maximizing the log-likelihood under multivariate normality. In high-dimensional settings, we estimate $\Lambda$ using principal component analysis (PCA) (Bai, 2003), which yields a row-consistent estimator for $\Lambda$, that is, each estimated row $\widehat{\Lambda}_{i\cdot}$ consistently estimates its true counterpart ${\Lambda}_{i\cdot}$. This approach does not require the covariance matrix $\text{Cov}(\epsilon_{X})$ to be diagonal.
Finally, we use cross-validation to select the tuning parameter $k$. Algorithm 2 summarizes our estimation procedure.
@@
1:Obtain $\widehat{q}$ from $\bf X$} (e.g., Onatski, 2010).
@@
5:Obtain $\widehat{SIV}={\bf X}B_{\widehat{\Lambda}^{\perp}}$.
[…]
@@ 4.2 Theoretical properties
In this section, we study the theoretical properties of the estimator $\widehat{\beta}$ in Algorithm 2. We consider two paradigms: (1) low-dimensional settings, where the dimension of exposure $p$ is fixed; and (2) high-dimensional settings, where $p$ grows with the sample size $n$. For the former, we show that under mild regularity conditions, $\widehat{\beta}$ is $\sqrt{n}$-consistent. For the latter, we show that under mild regularity conditions, $\widehat{\beta}$ achieves a Lasso-type error bound. We also demonstrate variable selection consistency in both scenarios. In our theoretical results, we do not require $\widehat{\Lambda}=\Lambda$; we only need certain norms of $\widehat{\Lambda}$ to be consistent with those of $\Lambda$, which can be achieved by classical estimators. We first introduce assumptions for the low-dimensional case.
@@
    All coefficients $\Lambda$}, $\beta$, and the function $g(\cdot)$ in models (1) and (3) are fixed and do not change as $n\to\infty$.
@@
    $U_{i}$, $\epsilon_{x,i}$, and $\epsilon_{y,i}$ are independent random draws from the joint distribution of $(U,\epsilon_{x},\epsilon_{y})$ such that $E(\epsilon_{x})=\bm{0}$, $E(U)=\bm{0}$, $\text{Cov}(\epsilon_{x})=D$, $\text{Cov}(U)=I_{q}$, and $(U,\epsilon_{x},\epsilon_{y})$ are mutually independent. Furthermore, assume that $\text{Var}(\epsilon_{y})=\sigma^{2}$ and $\max_{1\leq j\leq p}\text{Var}(X_{j})=\sigma_{x}^{2}$; these parameters are fixed and do not change as $n\to\infty$.
@@
[…]
Conditions B1B2 are standard assumptions for the low-dimensional setting. Given Condition A2, Condition B3 requires that the estimator for factor loadings is root-$n$ consistent. Condition B4 is the population version of the sparse eigenvalue condition (Raskutti et al., 2011, Assumption 3(b)).
@@
In Theorem 2, it is assumed that $\widehat{k}=s$. This is a standard condition in the $\ell_{0}$-optimization literature (e.g., Raskutti et al., 2011; Shen et al., 2013).
@@
    $sq^{2}\log(p)\log(n)/n\rightarrow 0$, $n=O(p)$, and $q+\log(p)\lesssim\sqrt{n}$, where $x\lesssim y$ means there exists a constant $C$ such that $x\leq Cy$.
@@
    Assume the random variables in models (1) and (3) satisfy $E(X)=\bm{0}$, $E(U)=\bm{0}$, and $\text{Cov}(U)=I_{q}$. We also assume $\epsilon_{y}$ is independent of $(X,U)$ and $\epsilon_{x}$ is independent of $U$. Furthermore, assume $\epsilon_{y}$, $\epsilon_{x,j}$, and $X_{j}$ are sub-Gaussian random variables with sub-Gaussian parameters $\sigma^{2}$, $\widetilde{\sigma}_{j}^{2}$, and $\sigma_{j}^{2}$, respectively. The parameters satisfy $\sigma^{2}\leq C_{4}$, and $C_{5}\leq\widetilde{\sigma}_{j}^{2},\sigma_{j}^{2}\leq C_{6}$ for some constants $C_{4},C_{5},C_{6}>0$.
@@
Condition C1 allows the number of exposures $p$ to grow exponentially with the sample size, while the number of latent confounders $q$ grows at a slower polynomial rate. Condition C2 is a standard assumption in high-dimensional factor analysis (Fan et al., 2013; Shen et al., 2016) for loading identification. Condition C3 assumes that the exposures $X_{j}$ are sub-Gaussian and that the noise level is bounded. Condition C4 is a standard assumption on minimum signal strength.
@@
Assume that Conditions A1A4 and C1C3 hold, and that the tuning parameter satisfies $\widehat{k}=s$. Then:
@@
where $b$ is the best linear predictor of $U$ from $X$. Their result depends on (i) $\|b\|_{2}=O\left({1}/{\sqrt{p}}\right)$ and (ii) the term $(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y}$ being independent of $X$ under their joint Gaussian assumption. In general, these conditions fail to hold under model (3), where $g(U)$ is an unknown function.